An investment costs $36,766.29 today. If it matures in 10 years, and the interest rate of 3.21% is compounded daily, how much will the security be worth at maturity?

Answer =

Cash Flow =

R =

N =
Answer #1

Security value at maturity   $ 50,681.82

Future value of security = Investment amount *(1+i)^n
= 36766.29*(1+0.0000879)^3650
= $ 50,681.82
Where,
i = 3.21%/365 =      0.0000879
n = 10*365 = 3650